suppose you have $12,500 to invest. Which investment yields the greater return after 5 years?
7.3% compounded monthly or 6.9% compounded quarterly?

Answer by mananth(16946)

Principal P = 12500
Amount= A
years=n 5.00
compounded 12 times a year t
Rate = 7.30 0.01
Amount = P*((n+r)/n)^n*t

Amount = = 12500 *(1+0.01)^5*12
Amount = 12500 *(1+0.01 )^60
12500 *(1.01)^60
Amount = 17986.53

Interest = 17986.53 - 12500 = $5,487
------
Principal P = 12500
Amount= A
years=n 5.00
compounded 4 times a year t
Rate = 6.90 0.02
Amount = P*((n+r)/n)^n*t

Amount = = 12500 *(1+0.02)^5*4
Amount = 12500 *(1+0.02)^20
12500 *( 1.02 )^20
Amount = 17598.03

Interest = 17598.03-12500 =$5,098
m.ananth@hotmail.ca